# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ 4.1 Hölder-DPO Is Provably Robust
$\operatorname*{\mathop{\rm argmin}}_{\theta}\underbrace{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})}[\log p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})]}_{\textrm{ignorable const.}}+\underbrace{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})}[-\log\sigma(g_{\theta}(\tilde{s}))]}_{\textrm{Original DPO loss}}=\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}D_{\mathrm{KL}}}[p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\,\|\,\sigma(g_{\theta}(\tilde{s}))].$ (6)
Thus, vanilla DPO can be understood as matching $\sigma(g_{\theta}(\tilde{s}))\approx p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})$ in terms of KL divergence. However, the KL divergence is well-known to be highly sensitive to data contamination [10, 31], and accordingly, DPO is not robust (Theorem 3). This fact motivates replacing the KL divergence with a more robust alternative. In this paper, we employ the Hölder divergence [52], a generalization of the KL divergence that encompasses several robust divergences, defined as follows:
@@
Let ϕ:+:italic-$\phi:\mathbb{R}_{+}\rightarrow\mathbb{R}$ be a continuous function such that $\phi(h)\geq-h^{1+\gamma}$ for all $h\geq 0$ and $\gamma>0$, and $\phi(1)=-1$. Then, the Hölder divergence is defined as
@@
where ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}S_{\gamma}(p\,\|\,q)}\coloneqq\phi\left(\frac{\mathbb{E}_{p}[q^{\gamma}(\omega)]}{\mathbb{E}_{q}[q^{\gamma}(\omega)]}\right)\cdot\mathbb{E}_{q}[q^{\gamma}(\omega)]$, $S_{\gamma}(p\,\|\,p)\coloneqq-\mathbb{E}_{p}[p^{\gamma}(\omega)]$, and $p$, $q$ are non-negative, non-zero functions.
@@
$\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}D_{\mathrm{H}}}[p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s}))\,\|\,\sigma(g_{\theta}(\tilde{s}))].$ (8)
@@
Prob. (8) generalizes the KL divergence and encompasses two robust divergences.
@@
$\gamma>0$ $-h^{1+\gamma}$ pseudo-spherical (PS) score [39, 38] $\gamma$-divergence [31, 52]
@@
$\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}S_{\gamma}}(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\,\|\,\sigma(g_{\theta}(\tilde{s}))).$ (9)
@@
Suppose that $\theta^{*}(\epsilon)=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(\tilde{p}^{(\epsilon)}_{\mathcal{D}}\,\|\,\sigma(g_{\theta}))$ and $\theta^{*}=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(p_{\mathcal{D}}\,\|\,\sigma(g_{\theta}))$. Let $\phi(h)$ be twice-differentiable, and let $0<\sigma(g_{\theta}(s))$ and $0<\gamma<\infty$. Then, the IF of the Hölder-DPO, excluding terms independent of $s_{\textrm{flip}}$, is given by:
$\mathrm{IF}_{\mathrm{H-DPO}}(z,\theta,p_{\mathcal{D}})\propto|\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})]|,$ (10)
@@
The full proof and the complete form of Eq. (10) are summarized in Appendices E.3 and E.4. This robustness result is expected, as both the DP divergence and the $\gamma$-divergence are well known to be robust to data contamination across various settings [10, 31, 37, 32]. The key difference between the IFs of DPO variants (Theorem 2) and our method (Theorem 4) lies in the term $\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}$. This subtle change induces diminishing IF weights, as shown in Figure 2, and ensures the redescending.
@@
$\approx\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}[{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}(1-\epsilon)}p_{\mathcal{D}}(s)\,\|\,\sigma(g_{\theta}(s))].$ (Corollary 1)
Thus, our Hölder-DPO objective can be interpreted as matching ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}(1-\epsilon)}p_{\mathcal{D}}(s)\approx\sigma(g_{\theta}(s))$. However, the scaling factor ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}(1-\epsilon)}$ remains on the target distribution side. The model extension idea [53] addresses this issue by introducing an additional scaling parameter ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi}$ into the model:
$\{\theta^{*}(\epsilon),\,\ {\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi^{*}}\}\approx\operatorname*{\mathop{\rm argmin}}_{\theta,\,{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi}}S_{\gamma}[{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}(1-\epsilon)}p_{\mathcal{D}}(s)\,\|\,{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi}\sigma(g_{\theta}(s))].$
Intuitively, if ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi\approx 1-\epsilon}$, the objective approximately reduces to the clean form: $S_{\gamma}(p_{\mathcal{D}}(s)\,\|\,\sigma(g_{\theta}(s)))$. In other words, the aligned model $\sigma(g_{\theta^{*}(\epsilon)}(s))$ approximates the distribution of the clean dataset, i.e., $p_{\mathcal{D}}(s)\approx\sigma(g_{\theta^{*}(\epsilon)}(s))$. Furthermore, the optimized scaling parameter ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi}$ provides an estimate of the contamination ratio ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\epsilon}$, using ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\epsilon\approx 1-\xi}$.
@@
For any fixed $\theta$, the optimal solution to the inner optimization problem $\operatorname*{\mathop{\rm argmin}}_{\xi}S_{\gamma}(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\,\|\,\xi\sigma(g_{\theta}(s)))$ is given in closed form by ${\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\xi^{*}}=\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}[\sigma(g_{\theta}(\tilde{s}))^{\gamma}]\,/\,\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}$. Then, the contamination ratio $\epsilon$ can be estimated by:
$\hat{\epsilon}=\min\{0,1-\hat{\xi}^{*}\},\quad\text{where}\quad\hat{\xi}^{*}\approx\frac{\frac{1}{N}\sum_{i=1}^{N}\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))^{\gamma}}{\sum_{i=1}^{N}\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))^{1+\gamma}},\,\,\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))\coloneqq\frac{\sigma(g_{\theta}(\tilde{s}^{(i)}))}{\sum_{i=1}^{N}\sigma(g_{\theta}(\tilde{s}^{(i)}))}.$ (11)
@@
$\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\widehat{S}_{\mathrm{DP}}}\left(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\,\|\,\xi\sigma(g_{\theta}(\tilde{s}))\right)\coloneqq\operatorname*{\mathop{\rm argmin}}_{\theta}-\frac{(1+\gamma)}{N}\sum_{i=1}^{N}\big{[}\sigma(g_{\theta}(\tilde{s}^{(i)}))^{\gamma}\big{]}+\frac{\gamma}{N}\sum_{i=1}^{N}\sigma(g_{\theta}(\tilde{s}^{(i)}))^{1+\gamma}.$ (12)
@@ 4.3 Algorithm to Detect Data Contamination
Algorithm 1 summarizes the overall procedure for detecting contaminated data points. As shown in Line 1, our proposed Hölder-DPO serves as a plug-and-play replacement for the standard DPO loss. Simply training with the objective in Eq. (12) yields a robust model alignment algorithm. For further analysis, Line 2 estimates the contamination ratio $\hat{\epsilon}$ without
@@
requiring access to a clean validation dataset—unlike prior work that depends on one [23, 56]. We then perform dataset valuation using the estimated clean-data likelihood $\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}^{(i)}))$, identifying the estimated mislabeled subset.
@@ 5 Experiments
We benchmarked Hölder-DPO against two robust DPO variants: R-DPO [23] and Dr. DPO [103]. We also compared to several heuristic baselines, including cDPO [73] and vanilla DPO [82]. As shown in Theorem 3, none of these baselines satisfy the redescending property, though they may offer empirical improvements in some settings. All experiments are conducted on a single NVIDIA A100 GPU (40 GB VRAM, 83.48 GB system RAM). For SFT of the reference model, we train for one epoch using context–response pairs $(x,y_{\text{win}})$. Each DPO variant is then trained for three epochs. We implement all methods using the Transformers [94, 102], TRL [96], and PyTorch [78] libraries. We set $\beta=0.1$ and $\gamma=2.0$, and all other hyperparameters follow the TRL defaults. Complete implementation details, including Hugging Face URLs to models and datasets used, are provided in Appendix H.

### v2

@@ 4.1 Hölder-DPO Is Provably Robust
$)].\displaystyle\operatorname*{\mathop{\rm argmin}}_{\theta}\underbrace{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})}[\log p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})]}_{\textrm{ignorable const.}}+\underbrace{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})}[-\log\sigma(g_{\theta}(\tilde{s}))]}_{\textrm{Original DPO loss}}=\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}D_{\mathrm{KL}}}[p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\,\|\,\sigma(g_{\theta}(\tilde{s}$))]. (6)
[…]
@@
Let $\phi:\mathbb{R}_{+}\rightarrow\mathbb{R}$ be a continuous function such that $\phi(h)\geq-h^{1+\gamma}$ for all $h\geq 0$ and $\gamma>0$, and $\phi(1)=-1$. Then, the Hölder divergence is defined as
@@
where ${\color[rgb]{1,0,1}S_{\gamma}(p\,\|\,q)}\coloneqq\phi\left(\frac{\mathbb{E}_{p}[q^{\gamma}(\omega)]}{\mathbb{E}_{q}[q^{\gamma}(\omega)]}\right)\cdot\mathbb{E}_{q}[q^{\gamma}(\omega)]$, $S_{\gamma}(p\,\|\,p)\coloneqq-\mathbb{E}_{p}[p^{\gamma}(\omega)]$, and $p$, $q$ are non-negative, non-zero functions.
@@
$)].\displaystyle\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}D_{\mathrm{H}}}[p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s}))\,\|\,\sigma(g_{\theta}(\tilde{s}$))]. (8)
@@
Prob. (8) generalises the KL divergence and encompasses two robust divergences.
@@
$\gamma>0$ $h^{1+\gamma}$ pseudo-spherical (PS) score [40, 39] $\gamma$-divergence [32, 53]
@@
[…]
@@
Suppose that $\theta^{*}(\epsilon)=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(\tilde{p}^{(\epsilon)}_{\mathcal{D}}\,\|\,\sigma(g_{\theta}))$ and $\theta^{*}=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(p_{\mathcal{D}}\,\|\,\sigma(g_{\theta}))$. Let $\phi(h)$ be twice-differentiable, and let $0<\sigma(g_{\theta}(s))$ and $0<\gamma<\infty$. Assume that $\phi(h)$ satisfies $\phi^{\prime}(h)\neq 0$ for $h>0$ and $\phi(h)\neq c\cdot h$ for any constant cc [^4] These assumptions are satisfied by the DP divergence and the PS score, as formally shown in Lemma 2. and the Hessian $\nabla_{\theta}^{2}\mathcal{L}(s,\pi_{\theta})|_{\theta=\theta^{*}}$ is positive definite. Then, the IF of the Hölder-DPO, excluding terms independent of $s_{\textrm{flip}}$, is given by:
$\mathrm{IF}_{\mathrm{H-DPO}}(s_{\textrm{flip}},\theta,p_{\mathcal{D}})\propto\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})],$ (10)
@@
The full proof and the complete form of Eq. (10) are summarised in Appendices E.3 and E.4. This robustness result is expected, as both the DP divergence and the $\gamma$-divergence are well known to be robust to data contamination across various settings [10, 32, 38, 33]. The key difference between the IFs of DPO variants (Theorem 2) and our method (Theorem 4) lies in the term $\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}$. This subtle change induces diminishing IF weights, as shown in Figure 2, and ensures the redescending.
@@
$\approx\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}[{\color[rgb]{1,0,1}(1-\epsilon)}p_{\mathcal{D}}(s)\,\|\,\sigma(g_{\theta}(s))].$ (Corollary 1)
Thus, our Hölder-DPO objective can be interpreted as matching ${\color[rgb]{1,0,1}(1-\epsilon)}p_{\mathcal{D}}(s)\approx\sigma(g_{\theta}(s))$. However, the scaling factor $\color[rgb]{1,0,1}(1-\epsilon)$} remains on the target distribution side. The model extension idea [54] addresses this issue by introducing an additional scaling parameter $\color[rgb]{1,0,1}\xi$} into the model:
$\{\theta^{*}(\epsilon),\,\ {\color[rgb]{1,0,1}\xi^{*}}\}\approx\operatorname*{\mathop{\rm argmin}}_{\theta,\,{\color[rgb]{1,0,1}\xi}}S_{\gamma}[{\color[rgb]{1,0,1}(1-\epsilon)}p_{\mathcal{D}}(s)\,\|\,{\color[rgb]{1,0,1}\xi}\sigma(g_{\theta}(s))].$
Intuitively, if $\color[rgb]{1,0,1}\xi\approx 1-\epsilon$}, the objective approximately reduces to the clean form: $S_{\gamma}(p_{\mathcal{D}}(s)\,\|\,\sigma(g_{\theta}(s)))$. In other words, the aligned model $\sigma(g_{\theta^{*}(\epsilon)}(s))$ approximates the distribution of the clean dataset, i.e., $p_{\mathcal{D}}(s)\approx\sigma(g_{\theta^{*}(\epsilon)}(s))$. Furthermore, the optimised scaling parameter $\color[rgb]{1,0,1}\xi$} provides an estimate of the contamination ratio $\color[rgb]{1,0,1}\epsilon$}, using $\color[rgb]{1,0,1}\epsilon\approx 1-\xi$}.
@@
For any fixed $\theta$, the optimal solution to the inner optimization problem $\operatorname*{\mathop{\rm argmin}}_{\xi}S_{\gamma}(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\,\|\,\xi\sigma(g_{\theta}(s)))$ is given in closed form by ${\color[rgb]{1,0,1}\xi^{*}}=\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}[\sigma(g_{\theta}(\tilde{s}))^{\gamma}]\,/\,\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}$. Then, the contamination ratio $\epsilon$ can be estimated by:
$\hat{\epsilon}=\min\{0,1-\hat{\xi}^{*}\},\qquad\text{where}\qquad\hat{\xi}^{*}:=\frac{\frac{1}{N}\sum_{i=1}^{N}\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))^{\gamma}}{\sum_{i=1}^{N}\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))^{1+\gamma}},\,\,\bar{\sigma}(g_{\theta}(\tilde{s}^{(i)}))\coloneqq\frac{\sigma(g_{\theta}(\tilde{s}^{(i)}))}{\sum_{i=1}^{N}\sigma(g_{\theta}(\tilde{s}^{(i)}))}.$ (11)
@@
$\operatorname*{\mathop{\rm argmin}}_{\theta}{\color[rgb]{1,0,1}\widehat{S}_{\mathrm{DP}}}\left(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\,\|\,\xi\sigma(g_{\theta}(\tilde{s}))\right)\coloneqq\operatorname*{\mathop{\rm argmin}}_{\theta}-\frac{(1+\gamma)}{N}\sum_{i=1}^{N}\big[\sigma(g_{\theta}(\tilde{s}^{(i)}))^{\gamma}\big]+\frac{\gamma}{N}\sum_{i=1}^{N}\sigma(g_{\theta}(\tilde{s}^{(i)}))^{1+\gamma}.$ (12)
@@ 4.3 Algorithm to Detect Data Contamination
Algorithm 1 summarises the overall procedure for detecting contaminated data points. As shown in Line 1, our proposed Hölder-DPO serves as a plug-and-play replacement for the standard DPO loss. Simply training with the objective in Eq. (12) yields a robust model alignment algorithm. For further analysis, Line 2 estimates the contamination ratio $\hat{\epsilon}$ without
@@
requiring access to a clean validation dataset—unlike prior work that depends on one [24, 58]. We then perform dataset valuation using the estimated clean-data likelihood $\sigma(g_{\theta^{*}(\epsilon)}(\tilde{s}^{(i)}))$, identifying the estimated mislabelled subset.
@@ 5 Experiments
We benchmarked Hölder-DPO against two robust DPO variants: R-DPO [24] and Dr. DPO [105]. We also compared to several heuristic baselines, including cDPO [76] and vanilla DPO [84]. As shown in Theorem 3, none of these baselines satisfy the redescending property, though they may offer empirical improvements in some settings. All experiments are conducted on a single NVIDIA A100 GPU (40 GB VRAM, 83.48 GB RAM). For SFT of the reference model, we train for one epoch using context–response pairs $(x,y_{\text{win}})$. Each DPO variant is then trained for five epochs. We implement all methods using the Transformers [96, 104], TRL [98], and PyTorch [80] libraries. We set $\beta=0.1$ and $\gamma=2.0$, and all other hyperparameters follow the TRL defaults. Complete implementation details, including Hugging Face URLs to models and datasets used, are provided in Appendix H.